# ASME Letters in Translational Robotics

### v1

@@
Robotic Wire Arc Additive Manufacturing (WAAM) is a metal additive manufacturing technology offering flexible 3D printing while ensuring high quality near-net-shape final parts. However, WAAM also suffers from geometric imprecision, especially for low-melting-point metal such as aluminum alloys. In this paper, we present a multi-robot framework for WAAM process monitoring and control. We consider a three-robot setup: a 6-dof welding robot, a 2-dof trunnion platform, and a 6-dof sensing robot with a wrist-mounted laser line scanner measuring the printed part height profile. The welding parameters, including the wire feed rate, are held constant based on the materials used, so the control input is the robot path speed. The measured output is the part height profile. The planning phase decomposes the target shape into slices of uniform height. During runtime, the sensing robot scans each printed layer, and the robot path speed for the next layer is adjusted based on the deviation from the desired profile. The adjustment is based on an identified model correlating the path speed to change in height. The control architecture coordinates the synchronous motion and data acquisition between all robots and sensors. Using a three-robot WAAM testbed, we demonstrate significant improvements of the closed loop scan-n-print approach over the current open loop result on both a flat wall and a more complex turbine blade shape.
@@ 1 INTRODUCTION
A robotic WAAM system typically involves a 6-dof welding robot combined with a 2-dof trunnion to enable alignment of metal deposition with gravity [10]. An addition 6-dof robot mounted with sensors provides comprehensive inspection coverage. Motion of these robotic devices need to be coordinated to ensure task execution and avoidance of constraints such as joint limits and collision. Planning techniques to avoid robot singularities [11] and selection of welding paths [12] are commonly employed. The standard approach involves preplanned paths with consideration of robot kinematics [13, 14]. Although many methods incorporate redundancy for added flexibility, effective strategies are needed to resolve it [15, 16]. To address these challenges, we consider a three-robot setup, including a 6-dof welding robot, a positioner with 2-dof trunnion platform and a 6-dof sensing robot with a wrist-mounted laser line scanning measuring the printed height profile. We use the motion planning approach for multiple robot arm from our previous work [17] for robot arms path planning and [3] gravity alignment.
@@
This paper presents a multi-robot framework for WAAM process monitoring and control. In the planning phase, the welding and inspection motion between the robots and the positioner are generated to meet the welding and inspection requirements. During the welding phase, we consider the robot torch path speed as the control input while holding the welding parameters, such as the wire feed rate at constant based on the materials used. The inspection robot measurements the layer height using a wrist-mounted laser scanner. The scans from the laser scanner are registered using the calibrated robot kinematics, followed by removal of outlier points to generate the height profile of the deposit. The torch speed for the next layer is then adjusted by the deviation from the desired profile. The adjustment is based on an identified model correlating the torch speed to the change in deposition height. The control architecture, described in [18], coordinates the motion and data acquisition between all robots and sensors. The closed loop WAAM framework has been applied in our testbed for printing a vertical wall, a complex turbine-blade-like geometry and a cylindrical geometry. The geometric fidelity of these parts outperform the open-loop baselines in both layer-by-layer and continuous-scan modes. We also demonstrate the repeatability of the manufacturing process, so that once the corrected motion profile is learned, the scanning step may be removed, thereby enhancing manufacturing efficiency. This paper is organized to present the scan-n-print framework in Section 2, the deposition model in Section 3, and experimental results and evaluation in Section 4. The source code and datasets related to this study are publicly available. [^1] https://github.com/rpiRobotics/Convergent_Manufacturing_WAAM/tree/main/scan_n_print
@@ 2.1 Overview
The proposed closed-loop WAAM framework is illustrated in Figure 2. After the deposition of each layer and scanning, the part geometry is reconstructed, with each scan registered to the positioner frame using the forward kinematics of the scanning robot and the positioner. Sampling the surface along the welding path gives the height profile of the top surface. For each layer, the slicing algorithm generate the desired layer height $\Delta h_{d}$. The measured height profile of the current layer is compared with the desired profile and the difference, together with the torch speed to height model, generates the robot motion for the next layer.
@@ 2.2 Motion Generation and Execution
Motion planning is required for both the WAAM system (6-dof welding robot and 2-dof trunnion) and the sensing robot. As described in [3], motion of the combined 8-dof system is determined by the torch position based on the slicing of the target geometry, the commanded path velocity, and the alignment requirement of the gravity direction with the previous layer. There is one redundant degree of freedom for the rolling motion along the torch. This redundancy resolved using resolved motion control subject to speed and joint limit constraints. The scanning motion of the sensing robot is designed to ensure complete coverage of the top layer. As illustrated in Figure 3, the scan path is along the known welding trajectory while maintaining the appropriate laser-scan standoff distance. To fully cover critical areas at the start and end of the weld, side scanning motion is also used. As for the welding motion, the scanning path is designed in the workpiece frame and then transformed to the positioner frame once the workpiece is placed on positioner. The robot motion is designed to ensure that the scans plane is perpendicular to the deposited layer. Since the gravity alignment is not needed, there is two redundant degrees of freedom (6-dof scanning robot and the 2-dof positioner). The same resolved motion control is used for redundancy resolution. Both the welding and scanning joint trajectories are segmented into uniform linear segments for robotic execution.
@@ 2.3 Deposition Height Profile
During run time, each layer is scanned immediately after printing. The laser scans form a 3D point cloud in the laser scanner frame. To transform the points to the robot frame, we used the robot joint angles and the inverse kinematics which requires robot kinematic calibration [19]. To mitigate noise, we apply statistical and cluster-based outlier removal. The cleaned point cloud enables extraction of the weld top profile by sampling along welding travel direction. With a well-defined point cloud representation of the weld, we extracting the layer deposition profile by slicing the point cloud along the printing direction. The height of each location is determined by the height of the points within the point cloud. Figure 4 displays the top layer height profile of the welded piece in the $x$-axis direction.
@@ 2.4 Motion Optimization
The motion optimization module aims to achieve uniform deposition height and correct welding defects. It is an essential step to prevent error accumulation and ensure consistent surface quality. We adjust only the torch path speed while maintaining constant wire feed rate to regulate the amount of deposited materials at each time. Denote relationship between the torch speed and the weld deposition height at a given feed rate $\Delta h=f(v)$, where $v$ is the torch speed and $\Delta h$ the deposition height. Empirical characterization of this model is described in Section 3. In general $f$ is a monotonically decreasing function as higher speed means less time for deposition and therefore smaller $\Delta h$.
@@
Figure 5 shows the measured heights and the corresponding commanding speeds for the motion segments along a typical layer.
@@ 2.5 Step-wise vs Continuous Scan-n-Print
The scan-n-print framework operates in step-wise or continuous modes. In the step-wise mode, robots alternate between welding and scanning, with the deposition profile processing and motion optimization occurring after the scanning motion. In the continuous mode, the scanning robot provides real-time feedback by looking ahead of the layer height, allowing on-the-fly processing and optimization. The continuous mode offers more immediate feedback and correction, and does not require turning the welding arc on and off (which compromises deposition uniformity) between layers. However, the implementation is more complex due to the coordinated motion between all three robots.
@@ 3 Welding Model
The identified coefficients are provided in Table 1. Note that the coefficients of the first-order terms are consistently very close to $-0.5$ which aligns with the equation derivation. We also showed the RMSE of the data after fitting the model.
@@ 4.1 Hardware Testbed
The hardware testbed is shown in Figure 1. It comprises a MA2010 welding robot with a D500B positioner, and a MA1440 inspection scanning robot, all from Yaskawa Motoman. Fronius 500i provides the welding power source for the cold metal transfer process. The Fronius weld controller is integrated with the robot controller DX200. The weld material is ER 4043, an aluminum alloy with lower melting point compared to materials like steel alloy or stainless steel. Its high thermal conductivity often leads to varying weld conditions and a higher occurrence of defects. For laser scanning, we use an MTI Pro-Track G Series laser scanner installed on the inspection robot. The scanner must maintain a standoff distance between 65mm to 125mm with a field of view of approximately 40mm to 60mm. The scans are collected at a frame rate of about 100 Hz.
@@ 4.2 Performance Comparison: Wall Structure
Abbildung 9: Open-loop baseline vs. Scan-n-Print for a Wall Structure
Figure 9 shows the printed pieces produced for the baseline and scan-n-print. Even through visual observation alone, it is evident that the quality of the printed piece has significantly better layer uniformity with the closed-loop approach. Note that defects tend to occur predominantly at the edges of the weld piece due to the arc on/off effects. Figure 10(a) shows the standard deviation of the layer height for both the open-loop baseline and closed-loop correction tests. In the open loop case, the height variation in each layer accumulates with the number of layers. To evaluate the print performance without the edge effects, we exclude the edge data. Figure 10(b) shows that the closed-loop approach still outperforms the open-loop approach. Table 2 shows the statistics and improvements of the height uniformity over all layers. For the wall geometry, the standard deviation of height improved 66%. Even without the edges where defects dominate, the performance improved by 60%.
@@
Tabelle 2: Mean height standard deviations across all layers, with and without the edges, of the WAAM wall, for open-loop baseline and scan-n-print correction cases.
@@
Abbildung 10: Height STD of each layer of the WAAM wall. Open-loop baseline vs closed-loop correction.
@@ 4.4 Continuous Scan-n-Print
Abbildung 15: Continuous Scan-n-Print of a cylinder.
@@ 4.5 Repeatability Test
Abbildung 16: Repeatability test.
@@
Abbildung 17: Repeatability test: height STD of each layer. Open-loop baseline vs closed-loop approach vs repeats

### v2

@@
Robotic Wire Arc Additive Manufacturing (WAAM) is a metal additive manufacturing technology offering flexible 3D printing while ensuring high-quality near-net-shape final parts. However, WAAM also suffers from geometric imprecision, especially for low-melting-point metal such as aluminum alloys. In this paper, we present a multi-robot framework for WAAM process monitoring and control. We consider a three-robot setup: a 6-dof welding robot, a 2-dof trunnion platform, and a 6-dof sensing robot with a wrist-mounted laser line scanner measuring the printed part height profile. The welding parameters, including the wire feed rate, are held constant based on the materials used, so the control input is the robot path speed. The measured output is the part height profile. The planning phase decomposes the target shape into slices of uniform height. During runtime, the sensing robot scans each printed layer, and the robot path speed for the next layer is adjusted based on the deviation from the desired profile. The adjustment is based on an identified model correlating the path speed to changes in height. The control architecture coordinates the synchronous motion and data acquisition between all robots and sensors. Using a three-robot WAAM testbed, we demonstrate significant improvements of the closed loop scan-n-print approach over the current open loop result on both a flat wall and a more complex turbine blade shape.
@@ 1 INTRODUCTION
A robotic WAAM system typically involves a 6-dof welding robot combined with a 2-dof trunnion to enable alignment of metal deposition with gravity [10]. An addition 6-dof robot mounted with sensors provides comprehensive inspection coverage. Motion of these robotic devices need to be coordinated to ensure task execution and avoidance of constraints such as joint limits and collision. Planning techniques to avoid robot singularities [11] and selection of welding paths [12] are commonly employed. The standard approach involves preplanned paths with consideration of robot kinematics [13, 14]. Although many methods incorporate redundancy for added flexibility, effective strategies are needed to resolve it [15, 16]. To address these challenges, we consider a three-robot setup, including a 6-dof welding robot, a positioner with 2-dof trunnion platform and a 6-dof sensing robot with a wrist-mounted laser line scanning measuring the printed height profile. We use the motion planning approach for multiple robot arm from our previous work [17, 18] for robot arms path planning and [3] gravity alignment.
@@
This paper presents a multi-robot framework for WAAM process monitoring and control. In the planning phase, the welding and inspection motion between the robots and the positioner are generated to meet the welding and inspection requirements. During the welding phase, we consider the robot torch path speed as the control input while holding the welding parameters, such as the wire feed rate at constant based on the materials used. The inspection robot measurements the layer height using a wrist-mounted laser scanner. The scans from the laser scanner are registered using the calibrated robot kinematics, followed by removal of outlier points to generate the height profile of the deposit. The torch speed for the next layer is then adjusted by the deviation from the desired profile. The adjustment is based on an identified model correlating the torch speed to the change in deposition height. The control architecture, described in [19], coordinates the motion and data acquisition between all robots and sensors. The closed loop WAAM framework has been applied in our testbed for printing a vertical wall, a complex turbine-blade-like geometry and a cylindrical geometry, using an aluminum alloy. The geometric fidelity of these parts outperform the open-loop baselines in both layer-by-layer and continuous-scan modes. We also apply our approach to a steel alloy, known for its higher melting point and greater thermal stability compared to aluminum alloys. The proposed techniques works well and shows improvements over the open loop printing. We demonstrate the repeatability of the manufacturing process, so that once the corrected motion profile is learned, the scanning step may be removed, thereby enhancing manufacturing efficiency. This paper is organized to present the scan-n-print framework in Section 2, the deposition model in Section 3, and experimental results and evaluation in Section 4. The source code, datasets and video related to this study are publicly available. [^1] https://github.com/rpiRobotics/Convergent_Manufacturing_WAAM/tree/main/scan_n_print  [^2] https://youtu.be/Ei2e2mBJZ5s
@@ 2.1 Overview
The proposed closed-loop WAAM framework is illustrated in Figure 2. After the deposition of each layer and scanning, the part geometry is reconstructed, with each scan registered to the positioner frame using the forward kinematics of the scanning robot and the positioner. Sampling the surface along the welding path gives the height profile of the top surface. For each layer, the slicing algorithm generates the desired layer height $\Delta h_{d}$. The measured height profile of the current layer is compared with the desired profile and the difference, together with the torch speed to height model, generates the robot motion for the next layer.
@@ 2.2 Motion Generation and Execution
Motion planning is required for both the WAAM system (6-dof welding robot and 2-dof trunnion) and the sensing robot. As described in [3], motion of the combined 8-dof system is determined by the torch position based on the slicing of the target geometry, the commanded path velocity, and the alignment requirement of the gravity direction with the previous layer. There is one redundant degree of freedom for the rolling motion along the torch. This redundancy resolved using resolved motion control subject to speed and joint limit constraints. The scanning motion of the sensing robot is designed to ensure complete coverage of the top layer. As illustrated in Figure 3, the scan path is along the known welding trajectory while maintaining the appropriate laser-scan standoff distance. To fully cover critical areas at the start and end of the weld, side scanning motion is also used. As for the welding motion, the scanning path is designed in the workpiece frame and then transformed to the positioner frame once the workpiece is placed on the positioner. The robot motion is designed to ensure that the scans plane is perpendicular to the deposited layer. Since the gravity alignment is not needed, there is two redundant degrees of freedom (6-dof scanning robot and the 2-dof positioner). The same resolved motion control is used for redundancy resolution.
@@
The overall system architecture to integrate multi-robot control and multi-sensor data acquisition is described in [3]. In the step-wise mode, the Cartesian motion of both the welding and scanning robots are segmented into uniform linear segments and encoded in the Motoman robot programming language, INFORM. The robot program together with the corresponding welding commands are sent to the Motoman DX200 robot controller. The DX200 controller coordinates the motion of the welding and scanning robots and trunnion table, and the Fronius weld controller. In the continuous mode, we bypass the Motoman controller and directly command the robot motion and weld controller. This is accomplished with two custom drivers using Robot Raconteur (RR) as the middleware [20]. The robot driver is developed using the Motoman MotoPlus package to stream robot motion setpoints and read robot joint angles at 8 ms to all three robots. The weld controller driver interfaces to the Fronius weld controller to change weld parameters and turn the welding arc on and off at 10 Hz.
@@ 2.3 Deposition Height Profile
During run time, each layer is scanned immediately after printing. The laser scans form a 3D point cloud in the laser scanner frame. To transform the points to the robot frame, we used the robot joint angles and the inverse kinematics which requires robot kinematic calibration [21]. To mitigate noise, we apply statistical and cluster-based outlier removal, and a average smoother along the welding direction. The cleaned point cloud enables extraction of the weld top profile by sampling along welding travel direction. With a well-defined point cloud representation of the weld, we extract the layer deposition profile by slicing the point cloud along the printing direction. The height of each location is determined by the height of the points within the point cloud. Figure 4 displays the top layer height profile of the welded piece in the $x$-axis direction.
@@ 2.4 Motion Optimization
The motion optimization module aims to achieve uniform deposition height and correct welding defects. It is an essential step to prevent error accumulation and ensure consistent surface quality. We adjust only the torch path speed while maintaining constant wire feed rate to regulate the amount of deposited materials at each time. Denote relationship between the torch speed and the weld deposition height at a given feed rate $\Delta h=f(v)$, where $v$ is the torch speed and $\Delta h$ the deposition height. Empirical characterization of this model is described in Section 3. In general, $f$ is a monotonically decreasing function as higher speed means less time for deposition and therefore smaller $\Delta h$.
@@
Figure 5 shows the measured heights and the corresponding commanding speeds for the motion segments along a typical layer. We apply a moving average smoothing filter to the height profile. This is a low pass filter that remove high frequency components in the profile that could cause unwanted high acceleration. The maximum and minimum speeds are also capped to prevent exaggerated movements. If a desired speed is not achieved because of the limited acceleration and speed, the remaining height errors are corrected in the subsequent layers. The optimized motion speed is commanded in the robot motion program if operated in the step-wise mode. In the continuous mode, we directly update the torch speed by adjusting the robot joint trajectory streamed to the robot controller in real time.
@@ 2.5 Step-wise vs Continuous Scan-n-Print
The scan-n-print framework operates in step-wise or continuous modes. In the step-wise mode, robots alternate between welding and scanning, with the deposition profile processing and motion optimization occurring after the scanning motion. In the continuous mode, the scanning robot provides real-time feedback by looking ahead of the layer height, allowing on-the-fly processing and optimization. The continuous mode offers more immediate feedback and correction, and does not require turning the welding arc on and off (which compromises deposition uniformity) between layers. However, the implementation of the continuous mode is more complex due to the coordinated motion between all three robots The motion optimization method is the same in both modes, but the update frequency is different. The step-wise mode optimizes the motion of the entire layer, while the continuous mode optimizes the motion of the current welding point during run-time using the look-ahead scanner. In our setup, the continuous mode closes the feedback loop at 100 Hz, limited by the scanning rate. Both modes produce high quality final parts. We observe that the continuous mode has slightly better performance in terms of the height standard deviation. This is because the step-wise mode requires turning the welding arc on and off at the edges.
@@ 3 Welding Model
The identified coefficients are provided in Table 1. Note that the coefficients of the first-order terms are consistently very close to $-0.5$ which aligns with the equation derivation. We also showed the RMSE of the data after fitting the model. During the scan-n-print operation, the wire feed rate is fixed at a chosen value and the torch speed is adjusted based on the scanning result.
@@ 4.1 Hardware Testbed
The hardware testbed is shown in Figure 1. It comprises a MA2010 welding robot with a D500B positioner, and a MA1440 inspection scanning robot, all from Yaskawa Motoman. Fronius 500i provides the welding power source for the cold metal transfer process. The Fronius weld controller is integrated with the robot controller DX200. The weld material is ER 4043, an aluminum alloy with lower melting point compared to materials like steel alloy or stainless steel. Its high thermal conductivity often leads to varying weld conditions and a higher occurrence of defects. In Section 4.4, we apply scan-n-print to ER 70S-6, a steel alloy with higher melting point and lower thermal conductivity than the aluminium alloy. For laser scanning, we use an MTI Pro-Track G Series laser scanner installed on the inspection robot. The scanner must maintain a standoff distance between 65mm to 125mm with a field of view of approximately 40mm to 60mm. The scans are collected at a frame rate of about 100 Hz.
@@ 4.2 Performance Comparison: Wall Structure
Abbildung 9: Open-loop baseline vs. Scan-n-Print for a Wall Structure, ER 4043 Aluminum Alloy.
Figure 9 shows the printed pieces produced for the baseline and scan-n-print. Even through visual observation alone, it is evident that the quality of the printed piece has significantly better layer uniformity with the closed-loop approach. Note that defects tend to occur predominantly at the edges of the weld piece due to the arc on/off effects. Figure 10(a) shows the standard deviation of the layer height for both the open-loop baseline and closed-loop correction tests. In the open loop case, the height variation in each layer accumulates with the number of layers. To evaluate the print performance without the edge effects, we exclude the edge data. We define the edges as within 7.5757.57.5mm from the welding arc on and off points. We can identify the regions on the point cloud using the collected robot joints data and the associated scans. Figure 10(b) shows that the closed-loop approach still outperforms the open-loop approach. Table 2 shows the statistics and improvements of the height uniformity over all layers. For the wall geometry, the standard deviation of height improved 66%. Even without the edges where defects dominate, the performance improved by 60%.
@@
Tabelle 2: Mean height standard deviations across all layers, with and without the edges, of the WAAM wall, for open-loop baseline and scan-n-print correction cases, ER 4043 Aluminum Alloy.
@@
Abbildung 10: Height STD of each layer of the WAAM wall. Open-loop baseline vs closed-loop correction, ER 4043 Aluminum Alloy.
@@ 4.3 Performance Comparison: Blade-like Structure
4.4 Performance Comparison: Steel Alloy
We follow the same procedure in Section 4.2 but using the steel alloy ER 70S-6. We select the best combination of wire feed rate and torch speed for the steel alloy: 200 ipm and 7 mm/sec. We apply the same model identification procedure in Section 3. The input height is approximately $\Delta h_{d}\approx 1.35$mm when the torch speed is set to 7mm/sec. The wall structure has the same dimension with width 65 mm and thickness 4 mm.
(a) Open-loop Baseline
(b) Scan-n-Print framework
Abbildung 15: Open-loop baseline vs. Scan-n-Print for a Wall Structure, ER 70S-6 Steel Alloy
Figure 15 shows the printed steel pieces for the baseline and scan-n-print. By visual observation, the open-loop welding with steel alloy does not create as much edge defects as with the aluminum alloy. This is because the steel alloy has higher melting point and lower thermal conductivity and is more stable than aluminum alloy. However, the top layer is still slightly curved compare to the close-loop approach. Figure 16(a) shows the standard deviation of the layer height for both the open-loop baseline and closed-loop correction tests. The open-loop case shows the error accumulation, but not as much as in the aluminum alloy case. The proposed scan-n-print still shows improvement. We exclude the edge data and evaluate the performance without edge effects. Figure 16(b) shows that the uniformity between the open-loop and the closed-loop approach is comparable. However, we see a trend of error accumulation in the open-loop case at the end of the weld and the height STD exceeds the closed-loop case significantly. Table 6 shows the statistics and improvements of the steel alloy case. As expected, the steel alloy is stable for WAAM so the mean height does not vary as much as the aluminum alloy in the open-loop baseline case (1.47 vs 0.74) but the proposed scan-n-print can further improve the weld by 26%. Without considering the edges, the performance improves by 10%.
Mean Height STD With edge Without edge
Baseline (mm) 0.74 0.31
Correction (mm) 0.55 0.28
Improvement 26% 10%
Tabelle 6: Mean height standard deviations across all layers, with and without the edges, of the WAAM wall, for open-loop baseline and scan-n-print correction cases, ER 70S-6 Steel Alloy.
(a) Full piece
(b) Edge excluded
Abbildung 16: Height STD of each layer. Open-loop baseline vs closed-loop correction, ER 70S-6 steel alloy.
@@ 4.5 Continuous Scan-n-Print
Abbildung 17: Continuous Scan-n-Print of a cylinder.
@@ 4.6 Repeatability Test
Abbildung 18: Repeatability test.
@@
Abbildung 19: Repeatability test: height STD of each layer. Open-loop baseline vs closed-loop approach vs repeats
@@ ACKNOWLEDGMENT
  • [18] He, H., Lu, C.-L., Saunders, G., Wason, J., Yang, P., Schoonover, J., Ajdelsztajn, L., Paternain, S., Julius, A., and Wen, J. T., 2024, “Fast and Accurate Relative Motion Tracking for Dual Industrial Robots,” IEEE Robotics and Automation Letters, 9(11), pp. 10153–10160.
@@
  • [20] Wason, J. D. and Wen, J. T., 2023, “Robot Raconteur®: Updates on an Open Source Interoperable Middleware for Robotics,” 2023 IEEE 19th International Conference on Automation Science and Engineering (CASE), Auckland, New Zealand, August 26 – 30, pp. 1–8, 10.1109/CASE56687.2023.10260569.